# Greedy Heuristics for Sampling-based Motion Planning in High-Dimensional State Spaces

### v1

@@ 2.3 State-space sampling
Prior work has been focused on increasing the efficiency of random sampling process of the sampling-based planning algorithms. In general, the performance of the planner highly depends on the distribution of random samples generated in each iteration. Many approaches rely on uniformly distributed samples over the whole state space to ensure that global optimality is preserved. However, the performance can often be improved by focusing the sampling on the regions with higher likelihood of finding better solutions.
@@
In Bialkowski et al. (2013), the authors propose a biased-sampling technique which generates random samples from the approximated obstacle-free distribution space. They show that the distribution converges to a uniform distribution over the free space as the number of samples approaches infinity. Similarly, Kim et al. (2014) decomposes the collision-free space into a set of spheres with varying radii called sampling cloud. Their Cloud RRT* algorithm exploits this sampling cloud to generate random samples and constantly update it to refine the current best solution. In P-RRT* and PQ-RRT* algorithms  (Qureshi and Ayaz, 2016; Li et al., 2020), random samples are directionally guided by artificial potential fields (Khatib, 1986) to efficiently explore the given environment. Unlike traditional rejection sampling approach (i.e., repeatedly sampling from the whole state space until a collision-free sample is found), these free-space biased sampling heuristic techniques quickly find better solutions and reduce the number of rejected samples however continue to sample the states that cannot improve the current solution.
@@
Prior work has been focused on improving the random sampling process and the convergence rate of RRT*. In Akgun and Stilman (2011), the authors proposed the bi-directional version of RRT*, which uses path-biased sampling and heuristic sample rejection techniques to accelerate the planner convergence rate in high-dimensional planning problems. Similarly, these path biasing and path refinement ideas have been employed in Islam et al. (2012); Alterovitz et al. (2011) to improve the original RRT* performance. However, these methods favor to find locally optimal solutions and decrease the likelihood of finding the samples that can improve the current solution. Moreover, techniques such as rectangular rejection sampling (Ferguson and Stentz, 2006; Otte and Correll, 2013) has also been applied to RRT* to further improve the convergence rate and reduce the computational requirements. But, these approaches do not increase the probability of finding better solutions and their effectiveness decreases as the number of state dimensions grows.
@@
To address these issues and accelerate the convergence of RRT*, direct informed sampling technique is introduced in Gammell et al. (2014, 2018), which is demonstrated with Informed RRT* algorithm. Informed RRT* behaves the same as RRT* until an initial solution is found. Once the initial solution is found, it biases the sampling on the subset of the state space (the $L^{2}$ informed set), which is an n-dimensional hyperellipsoid with a traverse diameter of current solution cost. Unlike aforementioned techniques, this approach maintains the uniform sample distribution over the informed subset and considers all the states that could provide better solutions, and is still effective in higher dimensional state spaces. However, the solution path found by the planner may contain redundant states, leading to produce tortuous paths with many twists and turns, especially in the early phase of the planning process. This can, in fact, result in an unnecessary path cost with large informed sets, slowing down the convergence rate of the planner. Therefore, recent works have a great interest in integrating more advanced path refinement techniques with informed sampling to further reduce the size of the hyperellipsoid, thereby accelerating the convergence (Kim and Song, 2015; Jiang et al., 2020; Wang et al., 2020b). Nevertheless, these advanced approaches can still be improved by incorporating greedy informed sampling heuristic (Section 3) to facilitate the exploitation process of the sampling-based planning algorithms.
@@
Research has also shown interest in integrating bio-inspired methods (Wang et al., 2018, 2019), learning sampling heuristics (Wang et al., 2020a; Zhang et al., 2021; Wang et al., 2022) to effectively guide the tree growth towards the goal by restricting the sampling state space of the problem domain, achieving non-uniform sampling distribution.
3 Greedy Informed Set
@@
Informed RRT* has been proposed to improve the convergence rate of the RRT* algorithm by utilizing direct informed sampling techniques, which performs sampling on the ellipsoidal subset of the planning domain, the $L^{2}$ informed set (Definition 6). This informed subset is an estimate of the omniscient set (Definition 5), the set of states that can provide a better solution, and both sets provide the theoretical upper bounds on the probability of improving a solution in problems seeking to minimize path length in $\mathbb{R}^{n}$—see Gammell et al. (2018) for more details. However, since the informed set is an admissible estimate of the omniscient set and approximated based on the current path cost, not all randomly sampled states on the informed subset may be able to lead to an improvement of the current solution path. In other words, this may result in a larger ellipsoidal informed set with redundant samples, taking unnecessary computation time for converging to the optimal solution, especially in higher state dimensions. For instance, considering the planning problems with a large number of obstacles and tight spaces, the initial solution paths produced by the algorithm can be tortuous, with many twists and turns, resulting in a substantial informed set with a higher solution cost. Therefore, for the purpose of improving the algorithm’s convergence rate and reducing the size of the subset in these constrained scenarios with tortuous solution paths, this paper proposes a new direct informed sampling strategy, which does not rely on the cost of the current solution path. It only uses the maximum heuristic cost in the existing solution, which could be obtained by finding the heuristic value of the greedily chosen state in the current solution path. This new informed subset only accepts randomly sampled states that could guarantee better solution costs than the heuristic value of the greedily chosen state (Definition 7).
@@
Figure 3: Illustration of the $L^{2}$ greedy informed set, $\mathcal{X}_{\hat{f}_{\text{greedy}}}$, for an $\mathbb{R}^{2}$ planning problem with path minimizing objectives. This greedy subset is highlighted in an ellipse with a dashed line and determined by the hypothetical minimum cost, $c_{min}$, from an initial state, $\mathbf{x}_{I}$, to the goal state, $\mathbf{x}_{G}$, and the heuristics cost of a state in the solution path with highest admissible estimates, $\hat{f}(\mathbf{x}_{max})$, as a transverse diameter.
@@
Definition 5 ((Omniscient set)).
Let $g(\mathbf{x})$ be the cost-to-come of a state, $\mathbf{x}\in\mathcal{X}$, the true cost of the optimal path from the initial state $\mathbf{x}_{I}$ to a state $\mathbf{x}$ and $h(\mathbf{x})$ be the cost-to-go of a state $\mathbf{x}$, the true cost of the optimal path from a state $\mathbf{x}$ to the goal state $\mathbf{x}_{G}$,
@@
Let $f(\mathbf{x})$ be the true cost of the optimal path from the initial state $\mathbf{x}_{I}$ to the goal state $\mathbf{x}_{G}$, constrained to pass through $\mathbf{x}$, $f(\mathbf{x})=g(\mathbf{x})+h(\mathbf{x})$. We then define the omniscient set, $\mathcal{X}_{f}$, as the set of states in collision-free space, $\mathcal{X}_{\text{free}}$, that can provide better solutions than the current solution cost, $c_{i}$,
@@
where .${\lVert.\rVert}_{2}$ denotes the $L^{2}$ norm (i.e., Euclidean distance) between two states. Let the function $\hat{f}(\mathbf{x})$ be a $L^{2}$ heuristic cost of a solution path from $\mathbf{x}_{I}$ to $\mathbf{x}_{G}$ constrained to pass through $\mathbf{x}$, i.e., $\hat{f}(\mathbf{x})=\hat{g}(\mathbf{x})+\hat{h}(\mathbf{x})$. We then define the $L^{2}$ informed set, $\mathcal{X}_{\hat{f}}$, as the set of states whose heuristic estimates can provide better solutions than the current solution cost, $c_{i}$,
@@
Moreover, this set is an upper bound estimate of the omniscient set and contains all the states that can yield better solutions, i.e., $\mathcal{X}_{\hat{f}}\supseteq\mathcal{X}_{f}$, if its heuristic estimates never overestimate the path’s true cost (admissible heuristics),
@@
Therefore, sampling or adding states from an admissible informed set is necessary to improve the current solution and maintain the uniform sample distribution over the subplanning problem for the asymptotic optimality guarantees (Gammell et al., 2018).
@@
However, in the case of complicated planning problems with many homotopy classes, the initial solution paths found by the planner may contain several twists and bends, resulting in larger informed sets. This can, in fact, slow down the convergence rate of the planner, especially in high state dimensions, since the heuristic does not provide substantial information; or in other words there is a very low chance of drawing a random sample from the informed set that could also belong to the omniscient set. Hence, instead of relying on the existing solution cost to construct these large informed sets, the states that found the existing solution path can be utilized to build up the decreasing series of ellipses that are bounded from above by the admissible informed sets (Definition 7).
Definition 7 (($L^{2}$ Greedy informed set)).
@@
Let $\mathbf{x}_{{max}}$ be the state in the current solution path, $\pi_{i}$, with maximum admissible heuristics,
$\mathbf{x}_{{max}}\coloneqq\underset{\mathbf{x}\in\pi_{i}}{\operatorname{\arg\max}}\left\{\hat{f}(\mathbf{x})\right\}.$ (10)
@@
We then define the $L^{2}$ greedy informed set, $\mathcal{X}_{\hat{f}_{\text{greedy}}}$, as the set of all the states in obstacle-free space that could produce better solution costs than the heuristic cost of this greedily chosen state $\mathbf{x}_{{max}}$ in the current solution (Figure 3),
$\mathcal{X}_{\hat{f}_{\text{greedy}}}=\left\{\mathbf{x}\in\mathcal{X}_{\text{free}}\mid\hat{f}(\mathbf{x})<\hat{f}(\mathbf{x}_{{max}})\right\}.$ (11)

### v2

@@ II-C State-space sampling
Prior work has focused on increasing the efficiency of the random sampling process in sampling-based planning algorithms. Generally, the performance of the planner depends significantly on the distribution of random samples generated in each iteration. Many approaches rely on uniformly distributed samples across the entire state space to ensure that global optimality is preserved. However, performance can often be improved by focusing the sampling on regions with a higher likelihood of finding better solutions.
@@
Bialkowski et al. [24] propose a biased-sampling technique that generates random samples from the approximated obstacle-free distribution space. They demonstrate that the distribution converges to a uniform distribution over the free space as the number of samples approaches infinity. Similarly, Kim et al. [25] decompose the collision-free space into a set of spheres with varying radii, called a sampling cloud. Their Cloud RRT* algorithm exploits this sampling cloud to generate random samples and continuously updates it to refine the current best solution. In the P-RRT* and PQ-RRT* algorithms [26, 27], random samples are directionally guided by artificial potential fields [28] to efficiently explore the given environment. Unlike the traditional rejection sampling approach (i.e., repeatedly sampling from the entire state space until a collision-free sample is found), these free-space biased sampling heuristic techniques quickly find better solutions and reduce the number of rejected samples; however, they continue to sample states that do not improve the current solution.
@@
Prior work has focused on improving the random sampling process and the convergence rate of RRT*. Akgun and Stilman [10] propose a bi-directional version of RRT* that uses path-biased sampling and heuristic sample rejection techniques to accelerate the planner’s convergence rate in high-dimensional planning problems. Similarly, these path biasing and path refinement ideas have been employed in [29, 30] to improve the performance of the original RRT*. However, these methods tend to favor finding locally optimal solutions, which decreases the likelihood of obtaining samples that can improve the current solution. Moreover, techniques such as rectangular rejection sampling [9, 11] have also been applied to RRT* to further improve the convergence rate and reduce computational requirements. However, these approaches do not increase the probability of finding better solutions, and their effectiveness decreases as the number of state dimensions grows.
@@
To address these issues and accelerate the convergence of RRT*, a direct informed sampling technique is introduced in [12, 31], which is demonstrated with the Informed RRT* algorithm. Informed RRT* behaves the same as RRT* until an initial solution is found. Once the initial solution is found, it biases the sampling on the subset of the state space (the $L^{2}$ informed set), which is an n-dimensional hyperellipsoid with a transverse diameter of the current solution cost. Unlike the aforementioned techniques, this approach maintains the uniform sample distribution over the informed subset and considers all the states that could provide better solutions and is still effective in higher-dimensional state spaces. However, the solution path found by the planner may contain redundant states, leading to tortuous paths with many twists and turns, especially in the early phases of the planning process. This can, in fact, result in an unnecessary path cost with large informed sets, thereby slowing down the convergence rate of the planner. Therefore, recent work has shown great interest in integrating more advanced path refinement techniques with informed sampling to further reduce the size of the hyperellipsoid, thereby accelerating convergence [32, 33, 34]. Nevertheless, these advanced approaches can still be improved by incorporating a greedy informed sampling heuristic (Section III) to facilitate the exploitation process of the sampling-based planning algorithms.
@@
Research has also shown interest in integrating bio-inspired methods [35, 36] and learning sampling heuristics [37, 38, 39] to effectively guide tree growth toward the goal by restricting the sampling state space of the problem domain, achieving a non-uniform sampling distribution.
III Greedy Informed Set
@@
Informed RRT* has been proposed to improve the convergence rate of the RRT* algorithm by utilizing direct informed sampling techniques, which perform sampling on the ellipsoidal subset of the planning domain, the $L^{2}$ informed set (Definition 6). This informed subset is an estimate of the omniscient set (Definition 5), the set of states that can provide a better solution, and both sets provide the theoretical upper bounds on the probability of improving a solution in problems seeking to minimize path length in $\mathbb{R}^{n}$—see [31] for more details. However, since the informed set is an admissible estimate of the omniscient set and is approximated based on the current path cost, not all randomly sampled states within the informed subset may lead to an improvement of the current solution path. In other words, this may result in a larger ellipsoidal informed set with redundant samples, taking unnecessary computation time to converge to the optimal solution, especially in higher-dimensional state spaces. For instance, in planning problems with a large number of obstacles and tight spaces, the initial solution paths produced by the algorithm can be tortuous, with many twists and turns, resulting in a substantial informed set with a higher solution cost. Therefore, to improve the algorithm’s convergence rate and reduce the size of the subset in these constrained scenarios with tortuous solution paths, this paper proposes a new direct informed sampling strategy that does not rely on the cost of the current solution path. Instead, it uses only the maximum heuristic cost from the existing solution, which can be obtained by finding the heuristic value of the greedily chosen state in the current solution path. This new informed subset accepts only randomly sampled states that can guarantee better solution costs than the heuristic value of the greedily chosen state (Definition 7).
@@
Figure 3: Illustration of the $L^{2}$ greedy informed set, $\mathcal{X}_{\hat{f}_{\text{greedy}}}$, for an $\mathbb{R}^{2}$ planning problem with path minimizing objectives. This greedy subset is highlighted in an ellipse with a dashed line and determined by the hypothetical minimum cost, $c_{\text{min}}$, from an initial state, $\mathbf{x}_{I}$, to the goal state, $\mathbf{x}_{G}$, and the heuristics cost of a state in the solution path with highest admissible estimates, $\hat{f}(\mathbf{x}_{\text{max}})$, as a transverse diameter.
@@
Definition 5 (Omniscient set).
Let $g(\mathbf{x})$ be the cost-to-come of a state, $\mathbf{x}\in\mathcal{X}$, which represents the true cost of the optimal path from the initial state $\mathbf{x}_{I}$ to a state $\mathbf{x}$, and let $h(\mathbf{x})$ be the cost-to-go of a state $\mathbf{x}$, which represents the true cost of the optimal path from a state $\mathbf{x}$ to the goal state $\mathbf{x}_{G}$,
@@
Let $f(\mathbf{x})$ be the true cost of the optimal path from the initial state $\mathbf{x}_{I}$ to the goal state $\mathbf{x}_{G}$, constrained to pass through $\mathbf{x}$, i.e., $f(\mathbf{x})=g(\mathbf{x})+h(\mathbf{x})$. We then define the omniscient set, $\mathcal{X}_{f}$, as the set of states in collision-free space, $\mathcal{X}_{\text{free}}$, that can provide better solutions than the current solution cost, $c_{i}$,
@@
where .${\lVert.\rVert}_{2}$ denotes the $L^{2}$ norm (i.e., Euclidean distance) between two states. Let the function $\hat{f}(\mathbf{x})$ be the $L^{2}$ heuristic cost of a solution path from $\mathbf{x}_{I}$ to $\mathbf{x}_{G}$ constrained to pass through $\mathbf{x}$, i.e., $\hat{f}(\mathbf{x})=\hat{g}(\mathbf{x})+\hat{h}(\mathbf{x})$. We then define the $L^{2}$ informed set, $\mathcal{X}_{\hat{f}}$, as the set of states whose heuristic estimates can provide better solutions than the current solution cost, $c_{i}$,
@@
Moreover, this set provides an upper bound estimate of the omniscient set and contains all the states that can yield better solutions, i.e., $\mathcal{X}_{\hat{f}}\supseteq\mathcal{X}_{f}$, if its heuristic estimates never overestimate the path’s true cost (admissible heuristics),
@@
Therefore, sampling or adding states from an admissible informed set is a necessary condition for improving the current solution and maintaining the uniform sample distribution to ensure asymptotic optimality guarantees [31].
@@
However, in the case of complex planning problems with many homotopy classes, the initial solution paths found by the planner may contain several twists and turns, resulting in larger informed sets. This can, in fact, slow down the convergence rate of the planner, especially in high-dimensional state spaces, since the heuristic does not provide substantial information. In other words, there is a very low chance of drawing a random sample from the informed set that could also belong to the omniscient set. Hence, instead of relying on the existing solution cost to construct these large informed sets, the states that found the existing solution path can be utilized to build a decreasing series of ellipses that are bounded from above by the admissible informed sets (Definition 7).
Definition 7 ($L^{2}$ Greedy informed set).
@@
Let $\mathbf{x}_{\text{max}}$ be the state in the current solution path, $\pi_{i}$, with the maximum admissible heuristic cost,
$\mathbf{x}_{\text{max}}\coloneqq\underset{\mathbf{x}\in\pi_{i}}{\operatorname{\arg\max}}\left\{\hat{f}(\mathbf{x})\right\}.$ (10)
@@
We then define the $L^{2}$ greedy informed set, $\mathcal{X}_{\hat{f}_{\text{greedy}}}$, as the set of all states in obstacle-free space that could produce better solution costs than the heuristic cost of this greedily chosen state $\mathbf{x}_{\text{max}}$ in the current solution (Figure 3),
$\mathcal{X}_{\hat{f}_{\text{greedy}}}=\left\{\mathbf{x}\in\mathcal{X}_{\text{free}}\mid\hat{f}(\mathbf{x})<\hat{f}(\mathbf{x}_{\text{max}})\right\}.$ (11)